# Prompting Large Language Models for Zero-shot Essay Scoring via Multi-trait Specialization

### v1

@@ 3.4 Comparisons with Other Methods
We compare MTS against an LLM-based zero-shot baseline as well as a supervised SOTA model.
Vanilla is an LLM-based zero-shot approach that asks LLM to assign a score directly within the target score range and provide evaluations before scoring in order to elicit CoT. Vanilla is in line with the "Require Rationale" approach proposed by Yancey et al. (2023) who showed that requesting rationales improves the performance in zero-shot setting. See Appendix C.2 for the full template.
NPCR (Xie et al., 2022) is the state-of-the-art supervised prompt-specific model in ASAP. It predicts the difference in the scores between an input essay and a reference essay. The predictions with multiple reference essays are averaged for inference. According to the result of our implementation (details in Appendix D), NPCR establishes a new state-of-the-art on TOEFL11 as well.
@@ 4 Main Results
The experimental results are summarized in Table 2. Overall, MTS consistently outperforms Vanilla in average QWK across all LLMs on both datasets, indicating that the advantage of MTS is universal across different types of LLMs. Moreover, the effectiveness of MTS remains solid across diverse settings of AES, given that ASAP and TOEFL11 represent distinctive environments in AES such as variations in essay genre and the first language (L1) backgrounds of the test takers.
The extent to which MTS improves over Vanilla in average QWK depends on different LLMs, varying from 0.023 to 0.355 on ASAP and from 0.212 to 0.437 on TOEFL11. Particularly, MTS greatly improves over Vanilla using Llama2-13b-chat, with gains of 0.355 ($0.205\rightarrow 0.560$) on ASAP and 0.329 ($0.238\rightarrow 0.567$) on TOEFL11. In general, we observe great performance gains for Llama 2 series and moderate gains for ChatGPT and Mistral-7b-Instruct. Interestingly, ChatGPT underperforms the small-sized Mistral-7b-Instruct on both datasets, implying that the model size might not be a decisive factor in the performance. Nevertheless, the larger model tends to perform better within the same model type, evidenced by the comparison between Llama2-7b-chat and Llama2-13b-chat.
Upon closer look, MTS surpasses Vanilla by a large margin in the majority of prompts, albeit with a few exceptions. For example, when utilizing Llama2-7b-chat, MTS achieves a QWK of 0.540 on Prompt 3 of TOEFL11, whereas Vanilla lags behind with 0.026, indicating a limited performance. This emphasizes the need for a well-designed prompt to evoke the LLM’s ample potential in essay scoring.
In terms of comparisons with the supervised SOTA, the gap in average QWK between MTS and NPCR can be narrowed down to 0.257 on ASAP and 0.171 on TOEFL11. This is promising given that NPCR consumes approximately a thousand labeled essays per prompt to reach the SOTA.
@@ 5.1 Analysis of Multi Trait Decomposition
Role of Rubric Guideline and Scoring Criteria.
When decomposing the writing proficiency into multiple traits, MTS (1) refers to rubric guidelines designed by human raters, and (2) generates scoring criteria along with each trait. We conduct ablation study over both options to investigate their roles in the model’s behavior.
As shown in Figure 4, the reference to the rubric guidelines has negligible impact on the average QWK. However, skipping the reference leads to higher standard deviation ($0.044\rightarrow 0.066$) of the QWKs, indicating an instable performance across the prompts. This could be attributed to the fact that skipping the reference yields generic and prompt-agnostic traits which does not entirely match the requirements articulated in the rubric guidelines of different prompts.
Next, the average QWK is declined greatly after discarding the scoring criteria, both with and without the reference to human standards. We conjecture that the scoring criteria regulates the LLMs’ behavior, encouraging them to adhere to the predefined criteria and preventing them from applying arbitrary scoring criteria to different essays.
@@ 5.2 Ablation Study of Trait Specialization
A characteristic these three strategies share in common is to decompose a complex problem of assigning an overall score into simpler subproblems and more specific tasks. We observe that this idea succeeds in zero-shot AES, in addition to the well-studied complex reasoning tasks (Zhou et al., 2022).
@@
(a) Average QWK under different scaling methods, measured on ASAP.
(b) Distributions (KDE) of $\hat{y}$, estimated on ASAP Prompt 5 using Llama2-7b-chat.
Figure 7: Comparison of different scaling methods.
@@ 5.3 Analysis of Trait Aggregation and Scaling
Effect of the Number of Traits.
MTS aggregates multiple trait scores $\{\hat{y}_{j}^{(i)}\}_{j=1}^{N_{T}}$ by averaging them and scaling via outlier clipping plus min-max scaling. We examine the effect of the number of traits by selecting all possible subsets of the original four traits with varying cardinality of $n\in\{2,3,4\}$ and evaluating the average QWK for each cardinality, as shown in Figure 6. We observe a clear tendency where a higher number of traits leads to elevated performance, suggesting that different traits are complementary to each other. In other words, MTS takes advantages of diversified assessment of writing proficiency.
Significance of Outlier Clipping and Min-max Scaling.
We conduct a comparative analysis of various scaling methods, illustrating the results in Figure 7. We consider a simple baseline of fixed scaling where the extreme values of the prediction interval (0 and 10 in our case) are used for min-max scaling. As shown in Figure 6(a), the simple fixed scaling mostly fails, especially with Llama2-7b-chat showing significantly degraded performance. Conversely, min-max scaling greatly outperforms fixed scaling in average QWK, and clipping outliers before min-max scaling further brings small but consistent improvements.
Furthermore, Figure 6(b) shows Kernel Density Estimation (KDE) of $\hat{y}$, providing insights into how the scaling mechanism of MTS works: (1) The LLM displays a bias toward predicting within a specific and concentrated interval, as shown in fixed scaling; (2) Min-max scaling effectively addresses the scoring bias by spreading its predictions across the target score range; (3) Outlier clipping further alleviates the distortion of the distribution caused by the outliers.
@@ 6 Related Work
Recent studies have investigated the use of LLM as a reference-free evaluation metric for natural language generation (NLG) tasks (Fu et al., 2023; Chiang and Lee, 2023). Zheng et al. (2023) revealed that strong LLMs (e.g., GPT-4) highly matched human preferences when evaluating the quality of AI assistant’s respones in open-ended QA tasks. Furthermore, LLMs achieved state-of-the-art agreement with human judgements in translation quality assessment (Kocmi and Federmann, 2023) and summarization task (Wang et al., 2023), verifying LLM evaluators’ scalability to a variety of NLG tasks.
@@ 7 Conclusion
We use a single NVIDIA A40 for each model inference including Llama2-7b-chat, Llama2-13b-chat and Mistral-7b-Instruct-v0.2. With only one sample in each batch, running MTS on ASAP test set takes approximately one GPU day for Llama2-7b-chat and Mistral-7b-Instruct-v0.2 and two GPU days for Llama2-13b-chat. Running time of MTS for ChatGPT through OpenAI API was similar to that of Llama2-13b-chat. However, the inference time may vary depending on the rate limits of different tiers.
@@ Appendix B Guidance from Multi Trait Decomposition
We show the guidance generated by multi trait decomposition. It consists of three parts: (1) trait, (2) description of the trait and (3) scoring criteria. MTS use (1) and (2) for its role prompt (system message), and utilize (1) and (3) for its conversation. Note that ASAP utilizes varying rubric guidelines across the eight prompts, resulting in distinct guidance for each prompt. Conversely, TOEFL11 employs the same rubric guidelines across the eight prompts, leading to the same guidance for all prompts. Due to its considerable length, we only display the guidance for Prompt 1 of ASAP and the guidance for all prompts of TOEFL11.
@@ C.1 Template for MTS
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "PERSON", "
{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits) [^2] In the original ASAP dataset, all named entities are marked in the format of ”@named entity”. We convert this format to ”{named entity}” in order to make the boundaries of the named entities more explicit.. Please do not penalize the essay because of the anonymizations.
@@ C.2 Template for Vanilla
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "PERSON", "
{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ C.3 Template for Scoring All Traits Sequentially
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "PERSON", "
{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ C.4 Template for Scoring Each Trait Independently
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "PERSON", "
{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ Appendix D Details of Re-implementation of NPCR
NPCR Xie et al. (2022) was originally implemented on ASAP. We re-implement it on TOEFL11 with minimal adjustments: the predicted scores are scaled to $[1,5]$ and mapped to low/medium/high with respect to the thresholds $[2.25,3.75]$, which is consistent with Section 3.3. In addition, the number of epochs is reduced from $80$ to $20$ so that the time for training and inference is kept in an acceptable range. We use the train, dev and test split provided by Tetreault et al. (2013). Other settings are identical to the original implementation.

### v2

@@ 3.4 Comparisons with Other Methods
We compare MTS against two LLM-based zero-shot baselines as well as a supervised SOTA model.
Rubric Scoring (Mizumoto and Eguchi, 2023) is a zero-shot approach designed for TOEFL11 which provides the LLM with human rubrics and asks it to assign an overall score within the target range.
Vanilla is a zero-shot approach that asks the LLM to assign an overall score within the target range and provide rationales before the score to elicit CoT. Requiring rationales has been shown to improve zero-shot scoring performance on the CEFR scale (Yancey et al., 2023). See Appendix C.2 for the full template.
NPCR (Xie et al., 2022) is the state-of-the-art supervised prompt-specific model which predicts the score difference of two input essays based on pairwise ranking objective. For a valid comparison, we re-implement NPCR to ensure it is evaluated on the same test set as MTS. See Appendix D for the training details.
@@ 4 Main Results
The zero-shot evaluation results are shown in Table 2. Building upon asking the LLM to assign a score, providing human rubrics (Mizumoto and Eguchi, 2023) and requiring rationales (Vanilla) yield similar average QWK on TOEFL11. MTS not only leverages the scoring rubrics but also retrieves evidence in trait-specific manner, consistently and significantly outperforming the baseline(s) in average QWK across all LLMs, on both datasets. In general, MTS achieves great performance gains over the baselines for Llama 2 series and moderate gains for ChatGPT and Mistral-7b-Instruct. For instance, MTS greatly improves over Vanilla using Llama2-13b-chat, with gains of 0.355 ($0.205\rightarrow 0.560$) on ASAP and 0.329 ($0.238\rightarrow 0.567$) on TOEFL11. This reassures the importance of careful prompt design to elicit LLM’s ample potential to perform AES. Moreover, the superiority of MTS remains solid across diverse settings of AES reflected in the datasets, including variations in essay genre and the first language (L1) backgrounds of the test takers.
By comparing different LLMs, we observe that Mistral-7b-instruct achieves the most competitive performance under different prompting methods overall, reaching average QWK of 0.550 on ASAP and 0.587 on TOEFL11. Interestingly, ChatGPT underperforms the small-sized Mistral-7b-Instruct across all methods but Rubric Scoring, implying that the model size might not be a decisive factor in the performance. Nevertheless, the larger model tends to perform better within the same model type, evidenced by the comparison between Llama2-7b-chat and Llama2-13b-chat.
In terms of the comparison with the supervised SOTA, the gap in average QWK between MTS and NPCR can be narrowed down to 0.171 on TOEFL11 and 0.242 on ASAP using Mistral-7b-instruct, as shown in Figure 4. This can be promising given that NPCR consumes approximately a thousand labeled essays per prompt to reach the SOTA. We expect further reduction in the gap with more powerful LLMs.
@@ 5.1 Analysis of Multi Trait Decomposition
Role of Rubric Guideline and Scoring Criteria.
When decomposing the writing proficiency into multiple traits, MTS (1) refers to rubric guidelines designed by human raters, and (2) generates scoring criteria along with each trait. We conduct ablation study over both options to investigate impact on the performance.
Figure 5 shows that the reference to the rubric guidelines has negligible impact on the average QWK. Yet skipping the reference leads to higher standard deviation ($0.044\rightarrow 0.066$) of the QWKs, indicating a fluctuating performance across the prompts. Next, the average QWK drops greatly after discarding the scoring criteria, both with and without the reference to human standards. One possible reason could be that the scoring criteria regulates the LLMs’ behavior, encouraging them to adhere to the predefined criteria for better consistency.
@@ 5.2 Ablation Study of Trait Specialization
A characteristic these three strategies share in common is to decompose a complex problem of assigning an overall score into simpler subproblems and more specific tasks. We observe that this idea succeeds in zero-shot AES, in addition to complex reasoning tasks (Zhou et al., 2022).
@@
ASAP TOEFL11
C L7 L13 M7 C L7 L13 M7
Vanilla+ 0.455 0.189 0.303 0.522 0.463 0.164 0.196 0.496
MTS 0.430 0.484 0.560 0.550 0.526 0.462 0.567 0.587
Table 4: Average QWK of the overall (Vanilla+) and diversified (MTS) assessment. C: ChatGPT; L7/13: Llama2-7/13b-chat; M7: Mistral-7b-instruct. Best QWK in each column is in bold.
@@ 5.3 Analysis of Trait Aggregation and Scaling
Merit of Diversified Assessment.
MTS averages multiple trait scores $\{\hat{y}_{j}^{(i)}\}_{j=1}^{N_{T}}$ and scales the averaged scores to the target score range. We examine whether aggregating more trait scores improves performance by selecting all subsets of the original four traits with varying cardinality of $n\in\{2,3,4\}$ and evaluating the average QWK for each cardinality, as shown in Figure 7. We observe a clear tendency where a higher number of traits leads to elevated performance, suggesting that different traits are complementary to each other. In other words, MTS takes advantages of diversified assessment of writing proficiency.
To further inspect if diversified scoring brings more benefit than the overall evaluation, we consider a new baseline called Vanilla+ which predicts the overall score in the same range as MTS (from 0 to 10) and applies the same scaling method as MTS [^2] Since the overall scores given by Vanilla+ lack diversity, applying outlier clipping may result in all scores being clipped to a single value for some prompts, in which case we set their QWKs to zero. . As shown in Table 4, the average of trait scores leads to better estimates of the writing quality than the overall score in most cases.
ASAP TOEFL11
C L7 L13 M7 C L7 L13 M7
fixed 0.350 0.254 0.477 0.520 0.357 0.071 0.445 0.385
minmax 0.405 0.477 0.553 0.529 0.526 0.438 0.420 0.499
+clipping 0.430 0.484 0.560 0.550 0.526 0.462 0.567 0.587
Table 5: Average QWK under different scaling methods. clipping: outlier clipping. Best QWK in each column is in bold.
Figure 8: Distributions (KDE) of $\hat{y}$, estimated on ASAP Prompt 5 using Llama2-7b-chat.
Significance of Min-max Scaling and Outlier Clipping.
We conduct a comparative analysis of various scaling methods. We consider a simple baseline of fixed scaling where $\hat{y}_{min}$ and $\hat{y}_{max}$ are fixed to 00 and $10$ during min-max scaling. As shown in Table 5, fixed scaling mostly fails, especially with Llama2-7b-chat showing significantly degraded performance. Conversely, min-max scaling greatly outperforms fixed scaling in average QWK for most cases, and clipping outliers further brings consistent and sometimes crucial improvements over min-max scaling.
Figure 8 shows Kernel Density Estimation (KDE) of $\hat{y}$, providing insights into how the scaling mechanism of MTS works: (1) The LLM displays a bias toward predicting within a specific and concentrated interval, as shown in fixed scaling; (2) Min-max scaling effectively addresses the scoring bias by spreading its predictions across the target score range; (3) Outlier clipping further alleviates the distortion of the distribution caused by the outliers, contributing to robust performance.
@@ 6 Related Work
Recent studies have investigated the use of LLM as a reference-free evaluation metric for natural language generation (NLG) tasks (Fu et al., 2023; Chiang and Lee, 2023). Zheng et al. (2023) showed that strong LLMs (e.g., GPT-4) highly matched human preferences when evaluating the quality of AI assistant’s respones in open-ended QA tasks. Furthermore, LLMs achieved state-of-the-art agreement with human judgements in translation quality assessment (Kocmi and Federmann, 2023) and summarization task (Wang et al., 2023), verifying LLM evaluators’ scalability to a variety of NLG tasks.
@@ 7 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (62076008) and the Key Project of Natural Science Foundation of China (61936012).
@@ Acknowledgments
We use a single NVIDIA A40 for each model inference including Llama2-7b-chat, Llama2-13b-chat and Mistral-7b-Instruct-v0.2. With only one sample in each batch, running MTS on ASAP test set takes approximately one GPU day for Llama2-7b-chat and Mistral-7b-Instruct-v0.2 and two GPU days for Llama2-13b-chat. Running time of MTS for ChatGPT through OpenAI API was similar to that of Llama2-13b-chat.
@@ Appendix B Guidance from Multi Trait Decomposition
We present guidance from multi-trait decomposition, comprising: (1) trait, (2) trait description, and (3) scoring criteria. MTS uses (1) and (2) for the role prompt (system message) and (1) and (3) for conversations. ASAP uses different rubric guidelines for each of its eight prompts, generating distinct guidance for each, while TOEFL11 applies the same rubric for all eight, resulting in identical guidance. Due to length, we show only the guidance for ASAP Prompt 1 and all prompts for TOEFL11.
@@ C.1 Template for MTS
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "
{PERSON}", "{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits) [^3] In the original ASAP dataset, all named entities are marked in the format of ”@named entity”. We convert this format to ”{named entity}” in order to make the boundaries of the named entities more explicit.. Please do not penalize the essay because of the anonymizations.
@@ C.2 Template for Vanilla
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "
{PERSON}", "{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ C.3 Template for Scoring All Traits Sequentially
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "
{PERSON}", "{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ C.4 Template for Scoring Each Trait Independently
I have made an effort to remove personally identifying information from the essays using the Named Entity Recognizer (NER). The relevant entities are identified in the text and then replaced with a string such as "
{PERSON}", "{ORGANIZATION}", "{LOCATION}", "{DATE}", "{TIME}", "{MONEY}", "{PERCENT}”, “{CAPS}” (any capitalized word) and “{NUM}” (any digits). Please do not penalize the essay because of the anonymizations.
@@ Appendix D Details of Re-implementation of NPCR
NPCR Xie et al. (2022) was originally implemented on ASAP. For our re-implementation on ASAP, while we leave out the same test set as MTS, the remaining data is randomly divided into train set and validation set by $4:1$. We re-implement NPCR on TOEFL11 as well with minimal adjustments: we use the train, dev and test split provided by Tetreault et al. (2013). The test set is identical to that of MTS. The predicted scores are scaled to $[1,5]$ and mapped to low/medium/high with respect to the thresholds $[2.25,3.75]$, which is consistent with Section 3.3.
For both datasets, the number of epochs is reduced from $80$ to $20$ so that the time for training and inference is kept in an acceptable range. Other settings are identical to the original implementation.